Circular No. 6511

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) D. Bockelée-Morvan, Observatoire de Paris; R. Moreno, Institut de

Radio Astronomie Millimétrique (IRAM); N. Biver, P. Colom, J. Crovisier, E. Gérard, E. Lellouch, and H. Rauer, Observatoire de Paris; W. Dent and J. K. Davies, Joint Astronomy Centre, Hilo; D. Despois, Observatoire de Bordeaux; and G. Paubert, IRAM, report observations of HNC, HCN, CO, CH_{3}OH, and H_{2}S with the IRAM 30-m telescope and of OH with the Nançay radio telescope on Nov. 8–10: “The J(1-0) line of HNC at 90.7 GHz was detected with a line area of 0.040 ± 0.008 K km s^{−1}, on the brightness-temperature scale. For comparison, the HCN J(1-0) line intensity is 0.90 ± 0.04 K km s^{−1}. The 157-GHz lines of methanol show a rotational temperature of 45 ± 2 K. The following production rates were measured: CO, 1.3 × 10^{29} molecules/s; HCN, 1.1 × 10^{27}; CH_{3}OH, 2.6 × 10^{28}; H_{2}S, 1.3 × 10^{28}; OH, 1.1 × 10^{30}.”

GX 5–1 M. van der Klis and R. Wijnands, Astronomical Institute ‘Anton Pan-

nekoek’, University of Amsterdam; E. Kuulkers, Oxford University; F. K. Lamb and D. Psaltis, University of Illinois at Urbana; S. Dieters, University of Alabama at Huntsville (UAH); J. van Paradijs, UAH and University of Amsterdam; W. H. G. Lewin, Massachusetts Institute of Technology; and B. Vaughan, California Institute of Technology, report: “RXTE observations of the low-mass x-ray binary (and Z source) GX 5–1 made on Nov. 6 and 16 show high-frequency quasiperiodic oscillations (QPO). Power-density spectra of the 10–50-keV flux show two QPO peaks in the range 325–895 Hz. The centroid frequency of the stronger of the two peaks increases from 567 ± 14 to 895 ± 12 Hz, its width decreases from 230 ± 40 to 50 ± 40 Hz, and its rms amplitude decreases from 6.7 ± 0.5 to 2.0 ± 0.3 percent (10–50 keV) from the left end of the horizontal branch to the top of the normal branch in the x-ray color-color diagram (i.e., for increasing mass-accretion rate). The weaker of the two peaks is only detected in part of this range, at a frequency that increases with accretion rate from 325 ± 9 to 448 ± 25 Hz. The frequency separation of the two peaks is consistent with being constant near 325 Hz. The peaks are at least 1.3 to 2.4 times weaker in the band 1–10 keV than in the band 10–50 keV. GX 5–1 is the second Z source (after Sco X-1; IAUC 6319) to show kHz QPO.”

1996 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green